OpenStudy (anonymous):

Solve. 5x2 – 20 = 0 x = 5 and x = 4 x = 4 and x = 2 x = –5 and x = 20 x = 2 and x = –2

Parth (parthkohli):

\( \color{Black}{\Rightarrow 5(x^2 - 4) = 0}\) Divide both sides by 5. \( \color{Black}{\Rightarrow x^2 - 4 = 0 }\) Now factor \(x^2 - 4\).

OpenStudy (anonymous):

5x^2-20=0 you can factor 5 out so we do that like this 5(x^2-4) = 0 5(x - 2 )(x + 2 ) = 0 x - 2 = 0 add 2 to each side to isolate x x=2 and for the other one x + 2 = 0 subtract two from each side to isolate x x= -2 therefore, x=2,-2